\begin{thebibliography}{13}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Abadi et~al.(2016)Abadi, Agarwal, Barham, et~al.]{Abadi2016}
Mart{\'\i}n Abadi, Ashish Agarwal, Paul Barham, et~al.
\newblock Tensorflow: Large-scale machine learning on heterogeneous distributed
  systems.
\newblock \emph{arXiv preprint arXiv:1603.04467}, 2016.

\bibitem[Boyd et~al.(2011)Boyd, Parikh, Chu, Peleato, and Eckstein]{boyd2011}
Stephen Boyd, Neal Parikh, Eric Chu, Borja Peleato, and Jonathan Eckstein.
\newblock Distributed optimization and statistical learning via the alternating
  direction method of multipliers.
\newblock \emph{Foundations and Trends{\textregistered} in Machine Learning},
  3\penalty0 (1):\penalty0 1--122, 2011.

\bibitem[Dai et~al.(2017)Dai, Yin, and Jha]{dai2017}
Xiaoliang Dai, Hongxu Yin, and Niraj~K Jha.
\newblock Nest: A neural network synthesis tool based on a grow-and-prune
  paradigm.
\newblock \emph{arXiv preprint arXiv:1711.02017}, 2017.

\bibitem[Han et~al.(2015)Han, Pool, Tran, and Dally]{han2015}
Song Han, Jeff Pool, John Tran, and William Dally.
\newblock Learning both weights and connections for efficient neural network.
\newblock In \emph{Advances in Neural Information Processing Systems (NIPS)},
  pp.\  1135--1143, 2015.

\bibitem[Han et~al.(2016)Han, Mao, and Dally]{DeepCompression}
Song Han, Huizi Mao, and William~J Dally.
\newblock Deep compression: Compressing deep neural networks with pruning,
  trained quantization and huffman coding.
\newblock \emph{International Conference on Learning Representations (ICLR)},
  2016.

\bibitem[Kiaee et~al.(2016)Kiaee, Gagn{\'e}, and Abbasi]{kiaee2016alternating}
Farkhondeh Kiaee, Christian Gagn{\'e}, and Mahdieh Abbasi.
\newblock Alternating direction method of multipliers for sparse convolutional
  neural networks.
\newblock \emph{arXiv preprint arXiv:1611.01590}, 2016.

\bibitem[Krizhevsky et~al.(2012)Krizhevsky, Sutskever, and
  Hinton]{krizhevsky2012}
Alex Krizhevsky, Ilya Sutskever, and Geoffrey~E Hinton.
\newblock Imagenet classification with deep convolutional neural networks.
\newblock In \emph{Advances in neural information processing systems}, pp.\
  1097--1105, 2012.

\bibitem[LeCun et~al.(1998)LeCun, Bottou, Bengio, and Haffner]{lecun1998}
Yann LeCun, L{\'e}on Bottou, Yoshua Bengio, and Patrick Haffner.
\newblock Gradient-based learning applied to document recognition.
\newblock \emph{Proceedings of the IEEE}, 86\penalty0 (11):\penalty0
  2278--2324, 1998.

\bibitem[Leng et~al.(2017)Leng, Li, Zhu, and Jin]{Leng2017}
Cong Leng, Hao Li, Shenghuo Zhu, and Rong Jin.
\newblock Extremely low bit neural network: Squeeze the last bit out with admm.
\newblock \emph{arXiv preprint arXiv:1707.09870}, 2017.

\bibitem[Liu et~al.(2013)Liu, Fardad, Masazade, and Varshney]{Liu2013}
Sijia Liu, Makan Fardad, Engin Masazade, and Pramod~K. Varshney.
\newblock On optimal periodic sensor scheduling for field estimation in
  wireless sensor networks.
\newblock In \emph{Global Conference on Signal and Information Processing
  (GlobalSIP)}, pp.\  137--140. IEEE, 2013.

\bibitem[Simonyan \& Zisserman(2014)Simonyan and Zisserman]{simonyan2014}
Karen Simonyan and Andrew Zisserman.
\newblock Very deep convolutional networks for large-scale image recognition.
\newblock \emph{arXiv preprint arXiv:1409.1556}, 2014.

\bibitem[Wen et~al.(2016)Wen, Wu, Wang, Chen, and Li]{wen2016}
Wei Wen, Chunpeng Wu, Yandan Wang, Yiran Chen, and Hai Li.
\newblock Learning structured sparsity in deep neural networks.
\newblock In \emph{Advances in Neural Information Processing Systems}, pp.\
  2074--2082, 2016.

\bibitem[Yang et~al.(2016)Yang, Chen, and Sze]{yang2016}
Tien-Ju Yang, Yu-Hsin Chen, and Vivienne Sze.
\newblock Designing energy-efficient convolutional neural networks using
  energy-aware pruning.
\newblock \emph{arXiv preprint arXiv:1611.05128}, 2016.

\end{thebibliography}

\documentclass{article} 
\usepackage{iclr2018_workshop,times}
\usepackage[colorlinks,
            linkcolor=black,
            anchorcolor=black,
            citecolor=black,
            urlcolor=black
            ]{hyperref}
\usepackage{url}	
\usepackage{amsmath}

\title{\hspace{0.2in} Systematic Weight Pruning of DNNs using \\ Alternating Direction Method of Multipliers}

\author{Tianyun Zhang, Shaokai Ye, Yipeng Zhang, Yanzhi Wang \& Makan Fardad\\ Department of Electrical Engineering and Computer Science\\
Syracuse University, \\ Syracuse, NY 13244, USA \\
\texttt{\{tzhan120,sye106,yzhan139,ywang393,makan\}@syr.edu} \\
}

\begin{document}

\maketitle

\begin{abstract}
We present a systematic weight pruning framework of deep neural networks (DNNs) using the alternating direction method of multipliers (ADMM). We first formulate the weight pruning problem of DNNs as a constrained nonconvex optimization problem, and then adopt the ADMM framework for systematic  weight pruning. We show that ADMM is highly suitable for weight pruning due to the computational efficiency it offers. We achieve a much higher compression ratio compared with prior work while maintaining the same test accuracy, together with a faster convergence rate. Our models are released at \url{https://github.com/KaiqiZhang/admm-pruning}.
\end{abstract}

\section{Introduction}

Despite the significant achievements enabled by DNNs, their large model size and computational requirements will add a significant burden to state-of-the-art computing systems \citep{krizhevsky2012,simonyan2014,DeepCompression}, especially for embedded and IoT systems. As a result, a number of prior works are dedicated to \emph{weight pruning methods} in order to simultaneously reduce the computation and model storage requirements of DNNs, with minor effect on the overall accuracy. 

A simple but effective method has been proposed in \citet{han2015}, which prunes the relatively less important weights and performs retraining for maintaining accuracy in an iterative manner. This method has been extended and generalized in multiple directions, including energy efficiency-aware pruning \citep{yang2016}, structure-preserved pruning using regularization methods \citep{wen2016}, and employing heuristics motivated by VLSI CAD \citep{dai2017}. 
While existing pruning methods achieve good model compression ratios, they are heuristic, lack theoretical guarantees on compression performance, and require time-consuming iterative retraining processes.

To mitigate these shortcomings, we present a systematic framework of model compression, by (i) formulating the weight pruning problem as a constrained nonconvex optimization problem, and (ii) adopting the \emph{alternating direction method of multipliers} (ADMM) \citep{boyd2011} for systematic  weight pruning. 
Upon convergence of ADMM, we remove the weights which are (close to) zero and retrain the network.
Our extensive numerical experiments indicate that ADMM works very well in practice and is highly suitable for weight pruning.
Overall, we achieve a model that has much fewer weights and less computation than previous weight pruning work while maintaining the same test accuracy as the model before pruning. The proposed method has a faster convergence rate compared with prior works. Our models are released at \url{https://github.com/KaiqiZhang/admm-pruning}.

\section{Problem Formulation and Proposed Framework}
\label{gen_inst}

Consider an $N$-layer DNN, where the collection of weights in the $i$-th layer is denoted by ${\bf{W}}_{i}$. In a convolutional layer the weights are organized in a four-dimension tensor and in a fully-connected layer they are organized in a two-dimension matrix \citep{Leng2017}. The loss function associated with the DNN is represented by $f({\bf{W}}_1,\dots,{\bf{W}}_N)$. In this paper, our objective is to prune the weights of the DNN and therefore we minimize the loss function subject to constraints on the cardinality of weights in each layer. Thus, our training process solves
\begin{equation}
\begin{aligned}
& \underset{\left \{{\bf{W}}_{i} \right \}}{\text{minimize}}
& & f({\bf{W}}_{1},\dots,{\bf{W}}_{N}),
& \text{subject to}
& & {\bf{W}}_{i}\in {\bf{S}}_{i}=\left \{{\bf{W}}\mid \mathrm{card}({\bf{W}})\le l_{i} \right \}, \; i=1,\dots,N,
\end{aligned}
\end{equation}
where $\mathrm{card}(\cdot)$ returns the number of nonzero elements of its matrix argument and $l_{i}$ is the desired number of weights in the $i$-th layer of the DNN after pruning. A prior work \cite{kiaee2016alternating} uses ADMM for DNN training with regularization in the objective function, which can result in sparsity as well. On the other hand, our method directly targets at sparsity with incorporating hard constraints on the weights, thereby resulting in a higher degree of sparsity.

It is clear that ${\bf{S}}_{1},\dots,{\bf{S}}_{N}$ are nonconvex sets, and it is in general difficult to solve optimization problems with nonconvex constraints. 
A recent paper of \citet{boyd2011}, however, demonstrates that ADMM can be utilized to solve nonconvex optimization problems in some special formats. The above problem can be equivalently rewritten in ADMM form as
\begin{equation*}
\begin{aligned}
& \underset{\left \{{\bf{W}}_{i} \right \}}{\text{minimize}}
& & f({\bf{W}}_{1},\dots,{\bf{W}}_{N})+\sum_{i=1}^{N} g_{i}({\bf{Z}}_{i}),
& \text{subject to}
& & {\bf{W}}_{i}={\bf{Z}}_{i},
\end{aligned}
\end{equation*}
where $g_{i}(\cdot)$ is the indicator function of ${\bf{S}}_{i}$ 
\begin{eqnarray*}g_{i}({\bf{Z}}_{i})=
\begin{cases}
 0 & \text { if } \mathrm{card}({\bf{Z}}_{i})\le l_{i}, \\ 
 +\infty & \text { otherwise. }
\end{cases}
\end{eqnarray*}

The augmented Lagrangian \citep{boyd2011} of the optimization problem is given by
{\small
\begin{equation*}
 L_{\rho} \big( \{{\bf{W}}_{i} \},\! \{{\bf{Z}}_{i} \},\! \{{\bf{\Lambda}}_{i} \} \big)
 =
 f({\bf{W}}_{1},\dots,{\bf{W}}_{N})
 +
 \sum_{i=1}^{N} g_{i}({\bf{Z}}_{i})
 +
 \sum_{i=1}^{N} \mathrm{tr}\left [{\bf{\Lambda}}_{i}^T({\bf{W}}_{i}-{\bf{Z}}_{i})  \right ]
 +
 \sum_{i=1}^{N} \frac{\rho_{i}}{2} \left \| {\bf{W}}_{i}-{\bf{Z}}_{i} \right \|_{F}^{2},
\end{equation*}
}%
where the matrices $\left \{{\bf{\Lambda}}_{1},\dots,{\bf{\Lambda}}_{N}\right \}$ are Lagrange multipliers, the positive scalars $\left \{\rho_{1},\dots,\rho_{N}\right \}$ are penalty parameters, $\mathrm{tr}(\cdot)$ denotes the trace, and $\left \| \cdot \right \|_{F}^{2}$ denotes the Frobenius norm. With the scaled dual variable ${\bf{U}}_{i}=(1/\rho_{i}){\bf{\Lambda}}_{i} $ the augmented Lagrangian can be equivalently expressed as
{\small
\begin{equation*}
L_{\rho} \big( \{{\bf{W}}_{i} \},\! \{{\bf{Z}}_{i} \},\! \{{\bf{U}}_{i} \} \big)
=
f({\bf{W}}_{1},\dots,{\bf{W}}_{N})
+
\sum_{i=1}^{N} g_{i}({\bf{Z}}_{i})
+
\sum_{i=1}^{N} \frac{\rho_{i}}{2} \left \| {\bf{W}}_{i} -{\bf{Z}}_{i}+{\bf{U}}_{i} \right \|_{F}^{2}
-
\sum_{i=1}^{N} \frac{\rho_{i}}{2} \left \|{\bf{U}}_{i} \right \|_{F}^{2}.
\end{equation*}
}%
The ADMM algorithm proceeds by repeating, for k = 0,1, \dots , the following steps \citep{boyd2011,Liu2013}:
\begin{align}
  \left \{{\bf{W}}_{i}^{k+1} \right \}&:=\mathop{\arg\min}_{\left \{{\bf{W}}_{i} \right \}} \ \ \ L_{\rho}(\left \{{\bf{W}}_{i} \right \},\left \{{\bf{Z}}_{i}^{k} \right \},\left \{{\bf{U}}_{i}^{k} \right \}) \\
  \left \{{\bf{Z}}_{i}^{k+1} \right \}&:=\mathop{\arg\min}_{\left \{{\bf{Z}}_{i} \right \}} \ \ \ L_{\rho}(\left \{{\bf{W}}_{i}^{k+1} \right \},\left \{{\bf{Z}}_{i} \right \},\left \{{\bf{U}}_{i}^{k} \right \})  \\
  {\bf{U}}_{i}^{k+1}&:={\bf{U}}_{i}^{k}+{\bf{W}}_{i}^{k+1}-{\bf{Z}}_{i}^{k+1},
\end{align}
until both of the following conditions are satisfied
\begin{equation*}
\left \| {\bf{W}}_{i}^{k+1}-{\bf{Z}}_{i}^{k+1} \right \|_{F}^{2} \ \le \epsilon_{i}, \ \ \left \| {\bf{Z}}_{i}^{k+1}-{\bf{Z}}_{i}^{k} \right \|_{F}^{2} \ \le \epsilon_{i}.
\end{equation*}

Problems (2) simplifies to
\begin{equation*}
\begin{aligned}
& \underset{\left \{{\bf{W}}_{i} \right \}}{\text{minimize}}
& & f({\bf{W}}_{1},\dots,{\bf{W}}_{N})+\sum_{i=1}^{N} \frac{\rho_{i}}{2} \left \| {\bf{W}}_{i}-{\bf{Z}}_{i}^{k}+{\bf{U}}_{i}^{k} \right \|_{F}^{2}, \\
\end{aligned}
\end{equation*}
where the first term is the loss function of the DNN, and the second term can be considered as a special $L_2$ regularization. 
Since the regularizer is a quadratic norm the complexity of minimizing the above loss function (for example, via gradient descent) is the same as the complexity of solving $\underset{\left \{{\bf{W}}_{i} \right \}} {\text{minimize}}\ 
f({\bf{W}}_{1},\dots,{\bf{W}}_{N})$.
On the other hand, problem (3) simplifies to
\begin{equation*}
\begin{aligned}
& \underset{\left \{{\bf{Z}}_{i} \right \}}{\text{minimize}}
& & \sum_{i=1}^{N} g_{i}({\bf{Z}}_{i})+\sum_{i=1}^{N} \frac{\rho_{i}}{2} \left \| {\bf{W}}_{i}^{k+1}-{\bf{Z}}_{i}+{\bf{U}}_{i}^{k} \right \|_{F}^{2}. \\
\end{aligned}
\end{equation*}
Since $g_{i}(\cdot)$ is the indicator function of ${\bf{S}}_{i}$ the solution of this problem is explicitly found to be \citep{boyd2011}
\begin{equation}
  {\bf{Z}}_{i}^{k+1} = {{\bf{\Pi}}_{{\bf{S}}_{i}}}({\bf{W}}_{i}^{k+1}+{\bf{U}}_{i}^{k}),
\end{equation}
where ${{\bf{\Pi}}_{{\bf{S}}_{i}}(\cdot)}$ denotes Euclidean projection onto the set ${\bf{S}}_{i}$. Note that ${\bf{S}}_{i}$ is a nonconvex set, and computing the projection onto a nonconvex set is a difficult problem in general. 
However, the special structure of ${\bf{S}}_{i}=\left \{{\bf{W}}\mid \mathrm{card}({\bf{W}})\le l_{i} \right \}$ allows us to express this Euclidean projection analytically. Namely, the solution of (5) is to keep the $l_{i}$ elements of ${\bf{W}}_{i}^{k+1}+{\bf{U}}_{i}^{k}$ with the largest magnitudes and set the rest to zero \citep{boyd2011}. Finally, we update the dual variable ${\bf{U}}_{i}$ according to (4). This constitutes one iteration of the ADMM algorithm. 

We observe that the proposed framework exhibits multiple major advantages in comparison with the heuristic weight pruning method of \citet{han2015}. 
Our proposed method achieves (i) higher convergence speed compared to iterative retraining, and (ii) higher compression ratio, as we demonstrate next.

\section{Experiments}
\label{headings}

We implement the network pruning method in Tensorflow \citep{Abadi2016}. We have tested weight pruning on the MNIST benchmark using the LeNet-300-100 and LeNet-5 model \citep{lecun1998}. In our experiments on these networks, our proposed ADMM method converges in approximately 20 iterations. 
As mentioned before, problem (2) can be solved by gradient descent. 
In experiments, we found that the number of steps we need for solving problem (2) by gradient descent is approximately 1/10 of the number of steps for training the original network. On the other hand, problem (3) and (4) are straightforward to carry out, thus their computational time can be ignored. Therefore, the total computation time of the ADMM algorithm is approximately equal to training the original network twice. While the solution of problem (1) should render $W_i$ with only $l_i$ nonzero elements, the solution we obtain through ADMM contains additional small nonzero entries. 
To deal with this issue, we keep the $l_{i}$ largest magnitude elements of $W_{i}$, set the rest to zero and no longer involve these elements in training (i.e., we prune these weights). We then retrain the network. 

Table 1 shows that our pruning reduces the number of weights by 22.9$\times$ on Lenet-300-100. Table 2 shows that our pruning reduces the number of weights by 40.2$\times$ on Lenet-5. Our pruning will not incur accuracy loss and can achieve a much higher compression ratio on these networks compared with the work of \citet{han2015}, which reduces parameters 12$\times$ on both networks.
Furthermore, on Lenet-5 we can reduce the number of weights by 10$\times$ in convolutional layers, which is also higher than the 8$\times$ in the work of \citet{han2015}.
Although on Lenet-5 the number of weights in convolutional layers is less than fully connected layers, the computation on Lenet-5 is dominated by its convolutional layers. This means that our pruning can reduce more computation compared with prior work.

\begin{table}
\centering
\caption{Weights pruning result on Lenet-300-100 network(without incurring accuracy loss)}
\label{my-label}
\begin{tabular}{cccc}
\hline
Layer & Weights & Weights after prune & Percentage of weights after prune \\ \hline
fc1 & 784$\times$300=235.2k & 9.41k & 4\% \\
fc2 & 300$\times$100=30k & 2.1k & 7\% \\
fc3 & 100$\times$10=1k & 0.12k & 12\% \\ \hline
\multicolumn{1}{l}{Total} & 266.2k & 11.6k & 4.37\% \\ \hline
\end{tabular}
\end{table}

\begin{table}
\centering
\caption{Weights pruning result on Lenet-5 network(without incurring accuracy loss)}
\label{my-label}
\begin{tabular}{cccc}
\hline
Layer & Weights & Weights after prune & Percentage of weights after prune \\ \hline
conv1 & 5$\times$5$\times$1$\times$20=0.5k & 0.1k & 20\% \\
conv2 & 5$\times$5$\times$20$\times$50=25k & 2.25k & 9\% \\
fc1 & 800$\times$500=400k & 8k & 2\% \\
fc2 & 500$\times$10=5k & 0.35k & 7\% \\ \hline
Total & 430.5k & 10.7k & 2.49\% \\ \hline
\end{tabular}
\end{table}

\subsubsection*{Acknowledgments}

Financial support from the National Science Foundation under awards CNS-1739748 and ECCS-1609916 is gratefully acknowledged.

\bibliography{iclr2018_workshop}
\bibliographystyle{iclr2018_workshop}

\end{document}